archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    15016 sbg g5 a9 Closed 2015 10 14 11098 Dedicated Cloud Maintenance Rbx2a FS 15015 Electrical Maintenance Planned 2015 10 14 11097 E mail Incident exchange FS 15014 Exchange Hosted 2013 Closed 2015 10 14 11096 Hosting Incident all Plan hostings FS 15013 filerz318 Closed 2015 10 14 11095 Hosting Modernization all Plan hostings FS 15012 mysql In progress 2015 10 14 11094 Dedicated Cloud Incident Rbx2a FS 15011 head 2020a leclercv2 ovh net Closed 2015 10 14 11093 Dedicated servers Incident Roubaix 2 FS 15010 head 316a storagenasha ovh net Closed 2015 10 14 11092 Network Incident Backbone US FS 15009 vac 3 Closed 2015 10 14 11091 Datacenters Modernization Strasbourg 1 FS 15006 Electrical wiring Planned 2015 10 14 11090 Dedicated Cloud Incident Rbx2a FS 15005 head 2019a leclercv2 ovh net Closed 2015 10 14 11089 VPS Incident Infrastructure FS 15004 VPS Cloud BHS Closed 2015 10 14 11088 VPS Incident Infrastructure FS 15003 VPS Cloud RBX Closed 2015 10 14 11087 Dedicated Cloud Incident Rbx2a FS 15002 head 2155 Closed 2015 10 14 11086 VPS Maintenance Backend Core FS 15001 VPS Classic 2014 RBX Closed 2015 10 14 11085 Network Incident RBX 4 FS 15000 Rack switch 51B11 Closed 2015 10 14 11084 Network Maintenance GRA 2 FS 14999 gra1 24ab n56 Planned 2015 10 14 11083 Network Incident Roubaix RBX 2 FS 14998 rbx s7 6k Closed 2015 10 14 11081 Network Modernization Roubaix FS 14996 rbx 50ab a70 vrack Planned 2015 10 14 11082 Network Modernization Roubaix RBX 4 FS 14997 rbx4 113a a70 vrack Planned 2015 10 14 11080 Network Incident Roubaix1 FS 14995 Rack switch 06C08 In progress 2015 10 13 11079 Network Incident Roubaix RBX 2 FS 14994 Rack switch 21D05 Closed 2015 10 13 11078 VPS Maintenance Infrastructure FS 14993 VPS

    Original URL path: http://status.ovh.es/?order=lastedit&project=0&tasks=&string=&type=&sev=&dev=&due=&cat=&status=&order2=lastedit&sort=asc&sort2=desc&date=&project=&sort=desc&order2=proj&sort2=asc (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    2011 08 23 1761 ADSL SDSL Incident Backend Core FS 5741 Switch on LIB59 Closed 2011 08 30 1783 ADSL SDSL Modernization Backend Core FS 5760 Backbone in Lille Closed 2011 09 07 1786 ADSL SDSL Modernization Backend Core FS 5763 Optical switching buckle Closed 2011 09 09 1801 ADSL SDSL Modernization Backend Core FS 5777 Optical pendulum migration Haubourdin HAU59 Closed 2011 09 12 1803 ADSL SDSL Modernization Backend Core FS 5778 Switching optical loop LQN LZI ASC TRI HEL Closed 2011 09 12 1804 ADSL SDSL Incident Backend Core FS 5779 WAS59 Closed 2011 09 13 1818 ADSL SDSL Incident Backend Core FS 5793 Authentication problem on the network of collect Closed 2011 09 18 1822 ADSL SDSL Maintenance Backend Core FS 5796 Optical loop LIB59 SEC59 FAC59 BOI59 BLE59 PIL59 Closed 2011 09 19 1827 ADSL SDSL Modernization Backend Core FS 5802 Updating DSLAM lib59 1 Closed 2011 09 21 1844 ADSL SDSL Modernization Backend Core FS 5818 updating DSLAM ble59 1 Closed 2011 09 24 1862 ADSL SDSL Maintenance Backend Core FS 5837 TRI59 HEL59 Closed 2011 09 29 1901 ADSL SDSL Maintenance Backend Core FS 5881 Lille loop1 LIB59 SEC59 FAC59 BOI59 BLE59 PIL59 Closed 2011 10 07 1925 ADSL SDSL Modernization Backend Core FS 5898 IP addresses change Closed 2011 10 21 1986 ADSL SDSL Maintenance Backend Core FS 5909 Lille loop 2 HEL59 TRI59 ASQ59 LZI59 LQN59 20 10 11 Closed 2011 10 21 1990 ADSL SDSL Incident Backend Core FS 5966 PPP connection Closed 2011 10 19 2000 ADSL SDSL Incident Backend Core FS 5976 NRA TES49 Closed 2011 10 21 2017 ADSL SDSL Maintenance Backend Core FS 5992 Lille loop 3 SAN59 WAS59 RON59 TOU59 RSA59 STM59 25 10 11 Closed 2011 10 28 2042 ADSL SDSL Maintenance Backend Core FS

    Original URL path: http://status.ovh.es/?order=prog&project=0&tasks=&string=&type=&sev=&dev=&due=&cat=&status=&order2=lastedit&sort=asc&sort2=desc&date=&project=&sort=desc&order2=proj&sort2=asc (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Autoconfiguration modem Attached to Project ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details The modem autoconfiguration service encounters problems Date Thursday 08 October 2015 17 34PM Reason for closing Done Additional comments about closing Service

    Original URL path: http://status.ovh.es/?do=details&id=11035 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details We lost the announcements SFR of BGP on the LNS Bordeaux Customers have automatically switched to the backup LNS Date Tuesday 29 September 2015 20 47PM Reason for closing Done Comments 1 Comment by OVH Monday 28

    Original URL path: http://status.ovh.es/?do=details&id=10922 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    The DSLAM men75 1 dsl is no longer responding to ping Comments 3 Comment by OVH Thursday 24 September 2015 15 31PM The site is isolated There is a fiber cut between men75 sites and th2 and between men75 sites and bgl93 A ticket has been generated by our fiber operator Comment by OVH Thursday 24 September 2015 15 32PM Ongoing incident in the PARIS MENILMONTANT sector We have operators

    Original URL path: http://status.ovh.es/?do=details&id=10893 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    investigating Date Thursday 03 September 2015 15 29PM Reason for closing Done Comments 4 Comment by OVH Wednesday 02 September 2015 15 42PM The problem was identified on the nexus mar 3 n5 Comment by OVH Wednesday 02 September 2015 15 44PM A shutdown no shutdown interface has corrected the problem We are monitoring the equipment Comment by OVH Wednesday 02 September 2015 19 06PM We are currently having a

    Original URL path: http://status.ovh.es/?do=details&id=10709 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details Our fibre operator is performing work on the fibre optic cable that links NRA SPR69 and MPL69 The loop is not secured but no impact is scheduled Date Thursday 20 August 2015 11 42AM Reason for closing

    Original URL path: http://status.ovh.es/?do=details&id=10562 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details Our fibre operator is performing work on the fibre optic cable that links NRA SPR69 and MPL69 The loop is not secured but no impact is scheduled Date Thursday 20 August 2015 11 39AM Reason for closing

    Original URL path: http://status.ovh.es/?do=details&id=10561 (2015-10-15)
    Open archived version from archive



  •